The company has a target D/E ratio of 0.45 which it intends to revert to as soon as possible, while its current D/E ratio is 0.50. Currently, the company has a beta of 1.5. The tax rate is 28%, the risk free rate 7% and the market risk premium, 6%. A very similar company recently issued bonds with a YTM of 10%. The company has R15 000 in total assets, R5000 in total liabilities with a book cost of 5% and has R10 000 in equity. The company currently has EBIT of R1000 which it expects to stay the same for the foreseeable future. R5000 will be raised, either by debt or equity. If debt is raised, the company expects to issue bonds at a market related YTM with a coupon rate of 10%. if the company chose to use debt financing, what would it’s WACC be?
1
Expert's answer
2020-08-24T19:05:09-0400

we know the formula of weighted average cost of capital ( WACC ):

WACC = ("\frac{E}{V}" x Re ) + [ ("\frac{D}{V}" X Rd ) (1-Tc) ]

where:

E = market value of equity =R 10000

D = market value of debit = R 5000

V = E+D = 10000 + 5000 = R15000

Re = rate of equity = ?

Rd = rate of debit = ?

Tc = 28% =0.28

first we calculate Re

from risk free rate 7% and market risk premium rate = 6% and beta = 1.5

so , cost of equity = risk free rate + beta ( market risk premium rate - risk free rate )

cost of equity = 7 + [1.5(6-7)]

Re = 7 +[ 1.5 (-1)]

Re = 7 - 1.5

Re = 5.5 %

Re = 0.055

Now we calculated Rd

from ebit = 1000 , debit raised = 5000 and coupan = 10% = 0.10

Re = "\frac{1000(0.10)}{5000}"

Re = 2%

Re = 0.02

Now find WACC

WACC = ("\frac{E}{V}" x Re ) + [ ("\frac{D}{V}" X Rd ) (1-Tc) ]

= ("\frac{10000}{15000}" x 0.055 ) + [ ("\frac{5000}{15000}" x 0.02 )(1-0.28)

= ( 0.036 ) + [ (0.006) (0.72)]

= 0.036 + 0.004

= 0.040

WACC = 4%